bill533:

Find x. A) 3 B) 3.75 C) 4.25 D) 5 E) 6.67

bill533:

1 attachment
bill533:

C

bill533:

@Skatercat @dude @Gdeinward

bill533:

or A

Gdeinward:

I dont know. The solution is on the tip of my tongue, but im unsure.

bill533:

i figured out

bill533:

This is known as the Intersecting Chord Theorem.... We have that 3 * 5 = 4 * x divide both sides by 4 15 / 4 = x = 3.75